OpenStudy (anonymous):

You travel 36 miles downstream in 4 hours when the current is 2 mph. In order to accomplish this , you must be paddling the boat at???

OpenStudy (anonymous):

7mph

OpenStudy (anonymous):

rate * time = distance Distance covered floating downstream without rowing in 4 hours 2*4=8 Distance covered rowing downstream with a 2mph current and rowing at 7mph. 9*4=36 Let r = rowing speed (r+2)*4=36 Solve for r. r=7mph

OpenStudy (anonymous):

how did you figure that out?

OpenStudy (anonymous):

Does my posting make any sense?

OpenStudy (anonymous):

yes..that makes sense.Thank you very much!

OpenStudy (anonymous):

Have a good day.